Given that the sound intensity 4.0 m from the sound source is 5.0 x 10βˆ’6wπ‘šβˆ’2.



What is the intensity at 2.0 m?

1
Expert's answer
2021-12-23T10:48:43-0500

Explanations & Calculations


  • For this, you need to use the formula for intensity at a given point from the source.
  • You can use it separately for the given 2 situations and find the intensity in the second situation.

"\qquad\qquad
\begin{aligned}
\small I&=\small \frac{P}{A}\\
\small I&=\small \frac{P}{4\pi r^2}\\
\small I& \approx\frac{1}{r^2}\\
\\
\small 5\times10^{-6}& \approx \small\frac{1}{4^2}\cdots(1)\\
\\
\small I_1 & \approx \small \frac{1}{2^2}\cdots(2)\\
\\
\small \frac{I_1}{5\times10^{-6}}&=\small \frac{4^2}{2^2}
\end{aligned}"


  • Once you subject "\small I_1" , you get the answer.